She estimates the following model as Rent = β0 + β1 Bedroom + β2 Bath + β3 Sqft + ε. The following ANOVA table shows a portion of the regression results. df SS MS F Regression 3 5,694,717 1,898,239 50.88 Residual 36 1,343,176 37,310 Total 39 7,037,893 Coefficients Standard Error t-stat p-value Intercept 300 84.0 3.57 0.0010 Bedroom 226 60.3 3.75 0.0006 Bath 89 55.9 1.59 0.1195 Sqft 0.2 0.09 2.22 0.0276 The coefficient of determination indicates that ________.
Answer #1

Answer

coefficient of determination = SSR/SST

using the given data information , we get SSR = 5694717 and SST = 7037893

this gives us

R^2 = 5694717/7037893

= 0.8092

Coefficient of determination is 80.92%, which indicates that 80.92% variation in the dependent variable can be explained by the independent variable